### v3

# Confidential Serverless Computing

Abstract
Although serverless computing offers compelling cost and deployment simplicity advantages, a significant challenge remains in securely managing sensitive data as it flows through the network of ephemeral function executions in serverless computing environments within untrusted clouds. While Confidential Virtual Machines (CVMs) offer a promising secure execution environment, their integration with serverless architectures currently faces fundamental limitations in key areas: security, performance, and resource efficiency.
We present Wallet, a lightweight confidential computing system for secure serverless deployments. By employing nested confidential execution and a decoupled guest OS within CVMs, Wallet runs each function in a minimal “trustlet”, significantly improving security through a reduced Trusted Computing Base (TCB). Furthermore, by leveraging a data-centric I/O architecture built upon a lightweight LibOS, Wallet optimizes network communication to address performance and resource efficiency challenges.
Our evaluation shows that compared to CVM-based deployments, Wallet has a 4.3 smaller TCB, improves end-to-end latency (15–93%), achieves higher function density (up to 907), and reduces inter-function communication (up to 27) and function chaining latency (16.7-30.2); thus, Wallet offers a practical system design for confidential serverless computing.
1 Introduction
Serverless computing is a rapidly evolving cloud paradigm offering low costs and simplified deployment [70]. As serverless applications increasingly manage sensitive data, particularly in AI/ML, strong security measures are of utmost importance [22, 19]. However, securely managing data within serverless computing is particularly challenging due to their ephemeral and distributed execution, where short-lived functions are chained through network communication [8, 152, 79].
To this end, Confidential Virtual Machines (CVMs) [77, 16, 17] provide a promising approach for the secure execution of serverless workloads in untrusted cloud environments [33], as they protect the confidentiality and integrity of entire VMs and do not require significant application modifications [76, 18].
However, due to strict serverless computing requirements, simply deploying serverless workloads directly within CVMs reveals critical inherent limitations (§ 3). In particular, despite its portability benefits, this approach exhibits non-ideal security properties, incurs prohibitive performance overheads, and leads to costly resource inefficiencies. In terms of security, CVMs rely on a full guest OS to host serverless applications, resulting in a bloated Trusted Computing Base (TCB) that expands the attack surface due to potential vulnerabilities in the untrusted OS (§ 3.1). From a performance perspective, CVMs suffer from long boot times—often exceeding function execution duration—due to strict security mechanisms (e.g., memory encryption) (§ 3.2). They also introduce significant I/O communication overheads, which are particularly detrimental in function chaining scenarios (§ 3.3) and require costly per-instance attestation, adding considerable latency to short-lived executions (§ 3.6). Finally, resource efficiency is severely impacted as CVM-based serverless frameworks struggle with function scheduling (§ 3.4), primarily due to consolidation challenges stemming from ineffective memory deduplication and hardware density constraints (§ 3.5).
Despite the advancements in virtualization architectures, no existing approach can fully resolve the combined challenges of security, performance, and resource efficiency in confidential serverless computing. Traditional VMs provide hardware-based isolation but incur substantial boot delays [103], high inter-VM communication costs [156], and maintain large TCBs because of their full virtualization stacks. Lightweight VMs (e.g., FireCracker [3], Kata [42]) optimize function density and startup latency [102], and reduce TCBs by trimming hypervisor and guest OS components [3]; however, they still suffer from significant I/O overheads and designs that complicate security verification [41, 140, 148]. Specialized OS architectures (e.g., LibOSes [136, 27, 88], Unikernels [87, 100, 111]) deliver near-instant startups by compiling applications with minimal kernel libraries but require developer efforts to craft application-specific images and sacrifice compatibility. CVMs (e.g., AMD SEV-SNP [9], Intel TDX [77], ARM CCA [17]), in turn, provide hardware-based security properties but introduce significant network communication overheads [92, 130, 112]. In summary, while each approach addresses individual aspects of the problem, a holistic solution that reconciles all these concerns remains an open challenge.
As a solution, we propose Wallet, a lightweight confidential computing system that enables secure serverless deployments in untrusted clouds by fundamentally rethinking how sensitive workloads are isolated and executed inside CVMs. In Wallet, each function runs in its own trustlet—a minimal serverless process that executes within a secure environment inside a CVM. With this design, client requests are securely routed through the untrusted guest OS, while the actual execution occurs in trustlets that encapsulate only the critical application code and the minimal system support required to serve these functions. This arrangement leverages two key mechanisms: nested confidential execution and a decoupled guest OS architecture. Inspired by nested virtualization [24] and ARM’s TrustZone model [18, 133], we partition the CVM space into smaller nested isolated regions. By confining sensitive code to trustlets, Wallet significantly reduces the overall TCB and minimizes the attack surface, thereby overcoming the limitations of hosting workloads on a monolithic guest OS.
To achieve high performance and fast function instantiation, Wallet adopts several techniques. First, it employs pre-initialized process templates (inspired by Android’s zygote model [37, 89]) to expedite the bootstrapping of trustlets. Complementing this, a dynamically loadable LibOS enables trustlets to achieve the minimal startup latency characteristic of unikernel-based systems [100, 124, 166] while retaining flexibility for diverse workloads. Moreover, Wallet introduces a data-centric I/O architecture that leverages high-density function co-location and replaces conventional CVM networking for inter-function communication, optimizing communication paths via secure data objects. It also incorporates differential attestation-an innovative technique that reuses pre-measured components to reduce attestation latency by focusing on only the mutable parts during invocation.
Collectively, these mechanisms enhance resource efficiency. As process templates and the LibOS are designed to share memory, and trustlets are ephemeral—occupying resources only during function execution—Wallet achieves a higher density of function deployments per CVM. This tight integration of performance optimizations with a thin TCB not only accelerates startup and execution but also allows for more efficient resource utilization, thereby supporting a larger number of concurrent function requests within the same memory footprint.
We implement Wallet on AMD SEV-SNP [9] with source code and evaluation setup to be publicly available. Our evaluation shows Wallet reduces TCB by 4.3 compared to CVMs while achieving lower end-to-end latency in real serverless applications for both cold (85.10-93.44%) and warm (14.97%) starts. In large-scale simulations using Azure Functions traces [165, 142], Wallet achieves significantly lower invocation latency than CVMs (i.e., 5ms vs. 489s at the 50th percentile, 1.5s vs. 881s at the 99th percentile). Further, performance microbenchmarks indicate that Wallet’s cold starts are faster than CVMs and 35% than traditional VMs, while warm starts show negligible latency (<10.3 ms). Function communication latency decreases by thanks to Wallet’s data-centric networking architecture, while function chaining performance improves by over CVMs. Attestation latency is also reduced as Wallet minimizes the measurement time (0.25 ms in warm starts). Lastly, Wallet has a smaller memory footprint than CVMs (up to ) for the same number of concurrent functions, rendering Wallet an ideal solution for secure serverless deployments in untrusted clouds.
2 Confidential Virtual Machines (CVMs)
As an alternative to existing cloud serverless virtualization architectures, CVMs [59, 107, 69] (e.g., AMD SEV-SNP [9], Intel TDX [77], ARM CCA [17]) provide a promising way to address escalating security requirements while preserving compatibility [112]. Precisely, CVMs protect the VM code, data, and memory from unauthorized access, including the hypervisor, via hardware-enforced isolation and encryption.
During a CVM launch, the VMM loads initial guest code, data, and state but cannot access the guest memory or modify its state, beyond this point. Guest firmware configures the private memory and performs necessary measurements. The guest kernel must be CVM-aware to interact with the trusted hardware. Host communication uses controlled, encrypted hypercalls. CVMs support remote attestation to verify their initial state and platform integrity. While providing strong internal security, they require secure protocols (e.g., networking) for data in transit/at rest to maintain their security guarantees.
Currently, Linux supports Intel TDX, AMD SEV-SNP, and ARM CCA (the hardware support of ARM CCA is absent as of now). In addition, Gramine-TDX [88] minimizes the guest TCB by employing a LibOS architecture and specializing the OS to a specific application on the Intel TDX platform.
CVM networking architecture. Since the hypervisor cannot directly access the CVM’s memory, CVMs use an unencrypted shared buffer (bounce buffer [130]) for network communication and employ encrypted protocols (e.g., TLS). Specifically, typical CVM networking involves [92]: (i) guest kernel encrypts and copies data to a bounce buffer, (ii) hypervisor transfers it to the destination CVM’s bounce buffer, and (iii) the destination CVM copies and decrypts the data for future consumption. This pipeline causes delays due to multiple copies, cryptographic operations, and heavy context switches [158, 112].
CVM partitioning. To improve intra-CVM isolation, CVMs offer a layered privilege model [1, 72, 110] beyond ring protections. This partitioning assigns each vCPU a privilege level defining its memory access rights and allowed instructions. The highest privilege level can intercept events from lower levels, enabling the provision of secure services (e.g., vTPM [71, 114]) without hypervisor dependence, and potentially shrinking the TCB by trusting only the most privileged components.
Specifically, AMD SEV-SNP’s Virtual Machine Privilege Levels (VMPL) [10] offer four levels (VMPL-0 to VMPL-3), with VMPL-0 being the most privileged. Higher levels can modify lower-level permissions. Intel’s TD-partitioning [72] enables complete nested virtualization within a CVM, allowing up to three nested VMs with an L1 VMM fully controlling the guests.
3 Motivation
| Baseline |
Runtime |
Total | ||||||||
1,115 | 903 | — | 36 | — | 2,054 | ||||||||
1,115 | 903 | 1,809 | 8,001 | 1,757 | 13,585 | ||||||||
1,115 | 903 | 3,177 | 744 | 1,757 | 7,696 | ||||||||
— | 903 | 3,177 | 744 | — | 4,824 | ||||||||
— | 332 | — | 780 | — | 1,112 |
Confidential computing [33] offers a promising way to secure serverless workloads in the untrusted cloud. However, CVM-based serverless computing faces six key challenges due to mismatches between CVM constraints and serverless requirements, analyzed below (experimental setup is described in § 8).
3.1 Issue #1: Large Trusted Computing Base
Minimizing the TCB is crucial for reducing the attack surface in confidential serverless computing. While CVMs eliminate hypervisor dependencies, CVM-hosted serverless stacks retain large TCBs containing monolithic legacy software with redundant components for backward compatibility [88].
Table 1 showcases that the CVM TCB size (4.8M LoC) is smaller than traditional VMs (7.7M LoC) but much larger than LibOS alternatives (e.g., Gramine LibOS: 2M LoC). Thus, CVM-hosted serverless stacks do not adequately minimize the TCB. The inclusion of a full guest OS increases complexity and exposes a large attack surface. In contrast, Wallet achieves approximately 4 TCB reduction compared to CVMs and 2 compared to Gramine LibOS without compromising security.
3.2 Issue #2: Slow Boot Times
Low startup latency is vital for serverless functions, but CVMs introduce inherent boot penalties [65, 112] due to three security mechanisms [14, 74]: (i) costly per-guest key generation and CVM memory pre-encryption, (ii) explicit guest state/memory initialization and boot measurement calculation for attestation, and (iii) synchronous signed report generation and external validation delaying function execution.
To this end, we measure the boot time of different virtualization architectures (Figure 1), which indicate that CVMs exhibit a boot time of 8.3 seconds, 2.2 slower than standard VMs (3.7s), primarily due to VMM setup, OS initialization, and encrypted memory configuration. These inevitable security-induced boot delays often exceed function lifetimes, making CVMs impractical for latency-sensitive serverless computing. While Wallet has overall the highest initialization time ( 12s), it is an one-time cost. In contrast, with an initialized Wallet, we can achieve faster cold starts (2.4 s) than traditional VMs (3.7 s), and a near-negligible warm start latency (<10.3 ms). Notably, this warm start achieves even lower start-up times compared to the LibOS ( 28 ms) on its own running without isolation on the host, thanks to its efficient CoW-based function execution.
3.3 Issue #3: High Networking Overheads
High-performance networking is crucial in distributed serverless applications with frequent data exchanges (e.g., function chaining). However, CVMs introduce significant networking overheads due to strict isolation and the requirement for additional encryption (e.g., TLS) to secure data in transit, which involves heavy context switches [30, 112].
To gauge these overheads, we measure the communication costs between two instances of the same virtualization model with varying message sizes (Figure 1). We observe that CVMs consistently exhibit higher latency compared to traditional VMs (9-20%) and lightweight solutions such as Kata containers and Gramine (up to 59).
We further examine the network performance of CVMs with various Linux configurations. Figure 1 shows iPerf [78] throughput (UDP, 1460 bytes payload, client runs on the host). “swotlb” means that bounce buffer is enabled, and “vhost” implies using the vhost optimization. The SNP VM always uses a bounce buffer. We observe that the SNP VM suffers from performance degradation (only 17% of the VM), and the guest-side halt polling [46] (“-hpoll”) mitigates the performance overhead at the cost of excessive CPU usage [112]. In contrast, Wallet achieves lower latency than CVMs (1.4-27) leveraging function co-location and following a data-centric networking model for secure, yet performant, communication.
3.4 Issue #4: Inefficient Scheduling
Efficient scheduling is a critical challenge [35, 168, 98] in confidential serverless computing. While optimized two-level (global/local) schedulers [144, 161, 64] favor warm starts, the limited CVMs per node [15, 125, 73, 75] can lead to more cold starts, increasing invocation latency. Further, traditional local schedulers, such as the Completely Fair Scheduler (CFS), are also ineffective [168] for short functions [67].
Specifically, confidential serverless functions in isolated environments cannot share resources, restricting the flexibility of resource-aware schedulers [144]. The limited pool of instances per CVM node forces balancing instance reuse, increasing cold start ratios [168, 98]. In addition, switching execution context between CVMs incurs significant overheads (e.g., TLB/cache flushing, memory validation) [28, 112, 169]. Our large-scale system simulation using sampled [154] Azure Functions traces [165, 142] (30 minutes, 4,000 functions, 4.1 million invocations) shows CVMs having higher scheduling latency across all percentiles compared to VMs and microVMs (Figure 2) under the same scheduling model.
Thus, we conclude that standard scheduling is suboptimal for CVM-based serverless. Wallet achieves lower scheduling latency with better tail latencies by sustaining more warm instances per node and using a run-to-completion model [168].
3.5 Issue #5: Impractical Consolidation
Efficiently packing multiple functions onto shared hardware resources is vital in serverless computing [3, 95], but also challenging for monolithic CVM-based serverless deployments for three fundamental reasons. First, CVM’s encrypted memory makes effective memory deduplication difficult, if not technically impossible [126, 15, 73, 75], as each memory is encrypted with a different key. Second, hardware limits the number of CVMs per node [15, 125, 73, 75], creating a hard ceiling on function density. Lastly, CVMs’ heavy-weight stacks cause higher per-function memory overhead, reducing achievable density considering a fixed hardware budget.
Figure 2 presents the memory usage of concurrent functions in a node. CVMs have the steepest increase, consuming 168 GB for 500 functions. The respective number for MicroVMs is 8.5 GB. Figure 2 further highlights the limit of CVMs per node imposed by the hardware (500 in our case).
Thus, CVM-based serverless struggles to consolidate functions efficiently. In contrast, Wallet can host multiple isolated functions within a CVM (up to 907 less memory than CVMs), exceeding the hardware-imposed density limit.
3.6 Issue #6: High Attestation Overheads
Remote attestation [106] establishes trust in a CVM, but its second-scale latency per CVM is excessive for ms-scale serverless functions. Specifically, attestation involves measuring software components and generating a signed report. To ensure function execution integrity, additional measurements of the function and input/output are necessary. This process can become costly, especially in function chaining scenarios.
Figure 2 shows that the measurement of a full guest OS kernel for a Linux CVM can take up to 1.1 s, significantly increasing the overall startup latency. While long-running workloads can amortize this cost, the short execution times of serverless functions make the repetitive attestation overhead prohibitive. Our Wallet prototype significantly reduces the measurement cost to 0.25 ms for warm starts by reusing already calculated measurements across the same functions.
3.7 Summary and Problem Statement
CVMs face critical limitations when employed in serverless settings. Their large TCB, slow boot times, I/O communication overhead, and costly attestation render them ill-suited for serverless computing. These constraints further lead to inefficient scheduling and low function density due to hardware limitations and suboptimal resource management.
To this end, there exists prior work that focuses on providing specialized solutions to tackle these challenges [88, 65]. Specifically, Gramine-TDX [88] provides a slim, security-first OS kernel, based on Gramine LibOS [153], to run unmodified applications in CVMs while having a minimal TCB. However, it is mainly suited for single-tenant architectures, where each serverless function requires its dedicated CVM, which conflicts with confidential serverless computing requirements.
Precisely, Gramine-TDX cannot achieve efficient function consolidation due to hardware CVM limits, requires per-function attestation, and forces inter-function communication through expensive CVM networking. Further, Gramine-TDX cannot leverage function co-location opportunities within the same CVM in a secure manner, as it operates with system-wide privileges and lacks the fine-grained privilege separation mechanisms necessary to safely isolate multiple co-located functions from each other and from the underlying software layers. These architectural constraints make Gramine-TDX unsuitable for the multi-tenant, high-density, and communication-intensive nature of serverless workloads.
Problem statement. To materialize confidential serverless computing, we have to answer a critical question: how can we execute lightweight, short-lived functions in the untrusted cloud in a secure and verifiable manner while complying with the performance and scalability requirements?
4 Overview
As a solution, we present Wallet, a lightweight confidential computing system for secure serverless deployments in untrusted cloud environments.
4.1 Threat Model
Wallet extends the CVM threat model [9, 30, 6, 127] by excluding the guest OS from its trust boundaries. We consider adversaries controlling the guest OS attempting to compromise serverless functions within trusted processes, the host OS attacking the CVM via VM-VMM interfaces (e.g., network I/O), or deploying malicious functions to access neighboring functions and leak data. Users trust function providers, but serverless and platform providers are potentially malicious.
The trusted monitor is Wallet’s only trusted software, enabling remote attestation. We assume that the platform hardware/firmware functions correctly and deployed functions do not intentionally leak data. Users must employ encrypted protocols (e.g., TLS) for communication. Wallet does not address denial-of-service, physical, and side-channel attacks.
4.2 System Architecture
Wallet builds on CVMs and uses hardware-enforced privilege partitioning (e.g., VMPL [10], TD-partitioning [72]) to create lightweight trustlets — minimal serverless processes that execute within a secure environment inside a CVM. Figure 3 presents Wallet’s key components: the trusted monitor, the trusted processes, the data objects and the zygote registry.
The trusted monitor is a small, privileged, and nested software component within Wallet’s TCB. It manages the instantiation and invocation of serverless functions within lightweight protection domains isolated from the guest OS. It performs security-critical tasks such as memory management, function deployment, and attestation.
Within the CVM, Wallet introduces the notion of trusted processes for hosting serverless functions. These processes possess enclave-like features: (i) isolation from the guest OS, (ii) trusted boot, and (iii) remote attestation capabilities. They are categorized into zygotes and trustlets. Zygotes serve as pre-initialized templates containing runtime environments (e.g., Python), while trustlets are lightweight instances, derived from zygotes, to execute functions in an isolated manner. To facilitate inter-function communication, Wallet provides the abstraction of data objects, managed by the trusted monitor ensuring isolation. This design favors reduced boot times, efficient memory use via copy-on-write [5, 53, 116, 52, 96], and optimized data exchange between co-located functions.
Meanwhile, the untrusted guest OS runs at a lower privilege level and mediates communication between the serverless orchestration framework and the trusted monitor. It lies out of the TCB and is responsible for retrieving zygote images and functions from the zygote registry and providing them to the monitor for validation, loading, and execution.
4.3 Life of a Request
[…]
4.4 Design Principles and Primitives
We now present the design principles and primitives of Wallet that address the six key problems, outlined in § 3.
#1: Nested confidential execution. As a countermeasure to the large TCB, the hardware-limited scalability and the costly inter-function communication of CVM-based serverless deployments, Wallet introduces nested confidential execution. Inspired by nested virtualization [24, 83, 32, 109, 164]—not directly applicable to CVMs due to host VMM support requirements [66, 4]—Wallet leverages intra-CVM isolation [1, 72, 110] to partition runtime components [104]. This allows the secure coexistence of untrusted components in lower-privilege compartments, while critical services operate in highly privileged domains, reducing the software TCB.
In this model, the trusted monitor governs the nested execution environments and operates at the highest privilege level, managing security-critical tasks while restricting untrusted components to lower levels via strict, hardware-enforced page-level access control [1, 72]. Further, it optimizes inter-function communication with data exchange via shared memory, realizing fast chaining for co-located functions.
Nested confidential execution also improves scalability without compromising security. It enables higher function density within a single CVM compared to conventional deployments, which are constrained by hardware limits on concurrent CVM instances per node [15, 125, 73, 75].
#2: Decoupled guest OS architecture. Conventional CVMs include the, often bloated [65], guest OS in their TCB. Wallet decouples the guest OS from trusted operations and removes its boot and measurement process from the function invocation path, mirroring ARM TrustZone’s principles, i.e., secure/normal world separation [18, 115], as performed in prior works, such as TLR (Trusted Language Runtime) [133].
After trusted monitor initialization, Wallet boots the untrusted guest OS, totally isolated from function execution environments. This decoupling excludes the guest OS kernel from the TCB, significantly reducing the attack surface and ensuring the guest OS vulnerabilities do not compromise the security of serverless functions. The guest OS handles only auxiliary tasks such as receiving requests and forwarding function invocations to the trusted monitor, maintaining compatibility with existing serverless frameworks. This reduces startup times compared to CVM-based deployments where cold starts involve full guest OS initialization and measurement.
#3: Trusted process templates. Existing confidential serverless deployments must initialize each function instance from scratch. Startup optimizations (e.g., snapshotting) are impractical [65] due to memory confidentiality (e.g., encryption) requirements. To this end, Wallet introduces trusted processes as the core abstraction for executing serverless functions.
These processes—categorized as zygotes and trustlets, inspired by Android’s process model [37, 89]—operate at a dedicated privilege level within isolated address spaces. Zygotes serve as pre-initialized templates with runtime dependencies, while trustlets instantiate from zygotes using copy-on-write mechanisms [5, 53, 116, 52, 96], enabling concurrent function launching with efficient memory usage and high density.
Building on this execution model, Wallet implements its differential attestation protocol, which incrementally builds cumulative trust chains by reusing zygote and function measurements, requiring only the measurement of mutable components (e.g., input/output) during invocation, thus reducing attestation latency for both standalone and chained functions.
#4: Dynamically loadable LibOS architecture. To minimize function startup latency and its TCB while supporting diverse workloads (e.g., Python, Node.js), Wallet employs a dynamically loadable LibOS architecture [166, 137], adopting concepts of unikernel-based systems [100, 124, 166]. However, instead of requiring developers to build custom LibOS images, Wallet dynamically loads lightweight LibOS-based runtimes tailored to specific functions at runtime. By providing the language runtimes that directly call a function, common serverless functions can be immediately used inside Wallet.
The LibOS forms the core of Wallet’s trusted processes, providing them with essential runtime functionalities. By dynamically loading minimal runtimes rather than full OSes or containers, Wallet reduces initialization overhead. This design allows on-demand instantiation of lightweight, pre-measured templates and avoids redundant software measurement steps during function invocation, further lowering startup latency without increasing developer effort.
#5: Confidential networking architecture. CVM I/O stacks [117, 44] suffer from performance penalties due to data copies, VM exits, and high CPU utilization for cryptographic operations [92, 112, 62]. Wallet’s key insight is to leverage its function density (§ 3.5) and provide a monitor-mediated, high-performance networking architecture for CVMs.
Wallet leverages function co-location and employs a data-centric I/O approach [161] to optimize function communication via secure data objects, building on insights from prior work (e.g., CAP-VM [135], Nephele [99], Pheromone [161]). Wallet’s high function density increases co-location opportunities, and thus, it can also benefit from data-centric schedulers [2, 82, 161] that effectively co-locate related functions.
In Wallet, data objects serve as secure endpoints for trustlets to exchange data without traversing conventional I/O stacks. The trusted monitor manages data object allocation and page table permissions, granting appropriate access to producers and consumers, and ensures secure data exchange while maintaining isolation between functions. This design eliminates unnecessary cryptographic operations and redundant data copies, reducing I/O overheads. For functions that cannot be co-located, Wallet falls back to standard CVM I/O, ensuring compatibility with existing serverless frameworks.
5 Design
5.1 Trusted Monitor
The trusted monitor is Wallet’s core component. It maintains strict isolation between each serverless function and the guest OS through hardware-enforced CVM partitioning and page table configurations. It operates at the highest privilege level of the CVM partitioning domains (e.g., VMPL-0).
Trusted process management. Wallet’s trusted monitor manages the CVM resources and preserves the state of the guest OS and trusted processes using process descriptors. A process descriptor includes the values for general-purpose and control registers, address space information (page tables), process ID, CVM-partitioning privilege level, object information (§ 5.4), and process state. The trusted monitor is responsible for allocating resources and scheduling the guest OS and the trusted processes. Internally, the descriptor state is linked to one vCPU state of the CVM, and the trusted monitor uses a special instruction to switch the context (e.g., AP creation call in SEV-SNP [14]). By executing the guest OS and the trusted processes at lower CVM partitioning levels, the trusted monitor governs their execution.
Isolation enforcement. The isolation between trusted processes and the guest OS requires careful design due to limited isolation domains (e.g., four domains for VMPL [10], three nested guests for TD-Partitioning [72]). To address this issue, the trusted monitor employs conventional ring protection, orthogonal to CVM partitioning: user functions run at ring3, monitor-managed code at ring0, preventing manipulation of the trusted processes’ page tables from ring3, and the guest OS at a lower privilege domain (i.e., VMPL-2). The page table-based isolation enables memory optimization through page sharing among functions—impossible in common CVMs—facilitating fast fork-based trustlet creation (§ 5.3.1) and efficient inter-function communication (§ 5.4).
Programming interface. As its execution environment base, Wallet adopts a LibOS architecture [23, 121, 31, 54, 91, 27]. This aligns well with the requirements of trusted processes, where the main code runs at the ring3, and offers users maximum flexibility through a common POSIX-compatible interface with a minimal resource footprint. Wallet tailors the LibOS architecture for serverless execution by allowing the dynamic loading of functions without requiring the recompilation of other components. Furthermore, the LibOS provides a data-centric I/O [161] API for inter-function communication (§ 5.2) and relies on function provider policies that include workflow information (e.g., function chaining), measurements of the involved zygotes and functions, and, optionally, remote storage encryption keys for functions requiring external storage access.
| Category | API | Description |
| Zygote | createZygote(void* image) -> zHandle | Loads and attests an image of a zygote and returns a zygote handle. |
| deleteZygote(zHandle) | Removes the zygote and its derived trustlets. | |
| Trustlet | createTrustlet(zHandle, fn) -> tHandle | Creates a trustlet based on a zygote and a provided function. |
| deleteTrustlet(tHandle) | Deletes the specified trustlet. | |
| invokeTrustlet(tHandle, void*) -> Result | Run a function in a trustlet and return the result. | |
| Attestation | attestMonitor() -> Report | Returns the attestation report of the monitor. |
| attest(Handle) -> Report | Retrieves the attestation report of a trusted process. | |
| Policy | loadPolicy(void*) | Loads the encrypted function provider policies. |
Monitor system calls. A serverless framework, running within the guest, interacts with the trusted monitor through the defined monitor system calls (Table 2). These calls facilitate the execution of serverless functions, including the creation of zygotes/trustlets, function invocation, performing attestation, and loading the provider’s policy. Additionally, the trusted monitor defines an interface between trusted processes for the LibOS to request backend operations (§ 5.2). All interfaces are kept small and concise to eliminate the risk of possible attack vectors introduced by internal oversight.
Attestation service. The trusted monitor includes an attestation service that generates an attestation report for remote users to verify the integrity of Wallet and its function execution. On function execution, the attestation service calculates the measurement of function components following Wallet’s differential attestation principles (§ 5.5) to optimize the measurement process by caching previously computed results. The root-of-trust of Wallet relies on CVM’s attestation mechanism [14, 74], whose attestation report includes the measurement of the initial guest state and memory. The attestation service integrates this report into its final result.
Performance consideration. The trusted monitor is a central part of Wallet’s operations. To ensure performant operations with multiple functions, the monitor allows each I/O operation to run simultaneously for multiple trustlets, without compromising its security properties. For memory accesses, the monitor is only called when (de)allocating memory. It manipulates the trustlets’ page table and sets the correct VMPL levels for the new pages, an operation which can also be parallelized.
5.2 LibOS Architecture
Wallet’s LibOS [23, 121, 31, 54, 91, 27] integrates a function runtime as well as a shim layer and filesystem with a nested namespace for efficient function execution.
Shim layer. The shim layer acts as the LibOS backend. While the LibOS provides most required OS services, certain operations mandate assistance from the trusted monitor and guest OS. The shim layer handles three key tasks: (i) memory allocation for trusted processes by coordinating with the trusted monitor, (ii) reading external host files, and (iii) handling data objects for function communication.
Filesystem. The LibOS provides an in-memory file system backed by a zygote, at the cost of a bloated image. Removing non-commonly-used files reduces image size but compromises serverless function flexibility and user experience. As a solution, Wallet employs a nested namespace [149] that integrates an embedded with an external filesystem. From the application’s perspective, all files appear at the same mount point (e.g., ‘/’). Internally, the LibOS first checks the embedded filesystem; if a file is missing, it retrieves it from the guest OS’s filesystem via the monitor. To ensure file integrity, the shim layer verifies files from the guest OS using provided measurements specified in a manifest [68, 153] embedded in the zygote image and aborts operations in case of a mismatch.
Data-centric I/O. For inter-function communication, the LibOS provides a data object abstraction to the trusted processes. This abstraction allows a function to create, get, and update data objects via the API shown in Table 3. Unlike conventional CVM networking that requires traversing entire I/O stacks with multiple context switches and encryption overhead [30, 92, 112, 62], Wallet employs a shared memory-based local object store [131, 161, 86, 122, 120, 29], enabling faster communication. This is especially effective for Wallet as it avoids the costly CVM network overhead (§ 3.3). When the serverless orchestrator fails to co-locate functions, the trusted monitor follows a hybrid I/O approach [101, 161] and resorts to the standard CVM I/O primitives via the guest OS, similar to the external file system access described above. Importantly, this mechanism ensures transparent communication between functions, regardless of whether they are co-located or distributed. § 5.4 details Wallet’s communication mechanism.
5.3 Trusted Process
A trusted process (Figure 5) is designed to efficiently host serverless functions within an isolated domain. Each trusted process encapsulates the LibOS, which is independent of the guest OS and provides functionalities to host a serverless runtime. Trusted processes run within the same CVM privilege domain, enabling memory sharing and CoW-based process creation. Meanwhile, the trusted monitor ensures their strict isolation by controlling the per-process page tables. Trusted processes are categorized into two types: zygotes and trustlets.
| API | Description |
| createObject(len,type) obj_id | Create data object. |
| getObject(obj_id) *obj | Get a pointer to the object obj_id. |
| getInputObject() (obj_id, len) | Get data object that contains input. |
| setOutputObject(obj_id) | Set data object as output. |
5.3.1 Zygote
A zygote is a template process containing a function execution environment tailored for a serverless runtime. Rather than running functions itself, it serves as a base for the instantiation of separate function execution contexts, namely trustlets (§ 5.3.2). They are fetched from the zygote registry, and after the successful integrity verification by the attestation service, the trusted monitor initializes their process descriptor and page tables. To optimize the function invocation process, a zygote offers a pre-initialization mechanism, enabling it to preload and initialize the function runtime environment [26]. During the zygote loading phase, the trusted monitor executes the zygote, and the zygote notifies about the completion of the initialization by making a specific monitor call. The trusted monitor then seals the initialized state and marks all zygote memory non-writable for the subsequent creation of trustlets.
5.3.2 Trustlet
A trustlet serves as the execution environment for serverless functions, inheriting the memory layout and execution environment from its base zygote. Each trustlet has data objects for inter-function communication. On trustlet creation, the trusted monitor duplicates the zygote’s process descriptor, loads the function, and inputs data to the corresponding memory region—all without copying the zygote memory.
Copy-on-write mechanism. Any trustlet write attempt to the non-writable zygote memory results in a page fault. To handle page faults, the trusted monitor installs a page fault handler for each trustlet. This handler delegates the actual page fault handling to the trusted monitor via a monitor call. Subsequently, the trusted monitor copies the page, updates the trustlet’s page table, and then resumes the execution.
Input and output. Function input and output are represented as special types of data objects. A function interacts with these objects via the API in Table 3. When invoking a trustlet, the trusted monitor creates an object for its input and retrieves the results from its output object after its execution.
Scheduling. The guest OS initiates the scheduling of trustlets via the invokeTrustlet() call. Wallet runtime enforces a run-to-completion execution model, i.e., Wallet executes a trustlet until completion for efficient execution [168], and returns the result to the guest OS. This fits for serverless workloads, which are most of the time short-lived. In case a function performs I/O tasks, the trustlet returns control back to the monitor and a different trustlet can be scheduled, even before the original one has finished its execution. On top of that, the scheduler is also responsible for handling the function chaining of the co-located trustlets (§ 5.4).
5.4 Confidential Network Architecture
[…]
Function chaining. Chaining multiple functions is a common pattern in serverless computing [36, 147, 162]. Wallet’s data object enables efficient function chaining through shared memory [131, 122]. Specifically, when chained functions are scheduled on the same host, the trusted monitor creates a special data object and grants write permission to the first function and read permission to the second function. If the functions can not be co-located, Wallet falls back to the normal network path: it returns the result of the first function to the serverless orchestrator, and then the orchestrator invokes the second function.
Note that, these chains are not limited to two functions and could potentially result in a circular chain, leading to the functions not producing a result. The current Wallet prototype does not consider this case. However, circular chains can be detected and prevented by the monitor in a future version. or Wallet’s policies could be extended to explicitly disallow it.
External network I/O. To facilitate network communications in Wallet across the CVM boundaries, additional steps are required. The monitor delegates the external network I/O handling to the guest OS. In practice, the monitor should encrypt the network data and provide it to the serverless framework for forwarding. Then, the serverless framework can send this data to the desired destination and get a response.
5.5 Differential Attestation
In confidential serverless computing, besides attesting the underlying platform, users need to verify the serverless runtime, the deployed function, and any inputs and outputs to validate the correct execution. This process can severely impact the function’s startup and execution latency. To alleviate this issue, Wallet introduces a differential attestation protocol to reduce the latency of verifying serverless function execution, which requires attesting the platform, runtime, function, and I/O. The key idea is to retain previously calculated measurements and compose a cumulative attestation report by measuring only the mutable parts at runtime for the function execution.
Security guarantees. Wallet’s differential attestation ensures the integrity of the monitor, trustlet, zygote, function, its input, and output. It relies on the correctness of Wallet, in particular on Wallet’s monitor, which is the only permanently immutable component of the system. All other parts of the differential attestation are at one point mutable, since they are fetched and instantiated by the monitor. After their initialization, all parts that constitute a trustlet become immutable with data protected with CoW. This enables the monitor to cache the hashes of these components for future use.
Attestation flow. Figure 7 illustrates the workflow of Wallet’s differential attestation protocol. It consists of two main phases: the initialization and the function invocation phase.
The initialization phase begins with the function provider sharing the public function key with the user, which is used to encrypt the user requests. Then, during the Wallet initialization phase, the function provider establishes a secure connection with the Wallet instance to share the function’s private key and the function’s policy that specifies the measurements of zygotes and functions permitted for use. The establishment of the secure connection relies on the attestation of the trusted monitor, whose trust is rooted in the ASP [14].
Precisely, the function provider first sends a nonce to the trusted monitor to prevent replay attacks. In turn, the attestation service generates a Diffie–Hellman (DH) key pair and retrieves the attestation report from the ASP, which includes the measurement of the trusted monitor, the provided nonce and the hash value of the DH public key as user data. The function provider receives and verifies the report and then sends the function private key and the policy to the Wallet instance. Wallet uses the function’s private key to decrypt user requests and the policy to attest the trusted processes.
Importantly, before providing private or security-sensitive data as a function input, a user must first perform a test function call, retrieve the attestation report, and verify that the Wallet instance is at the expected state and the correct toolchain is present. After this point, the user has gained trust in Wallet and can proceed with their function invocations.
The function invocation phase includes two scenarios depending on the Wallet’s state, namely cold and warm start (Figure 4). In the cold start case, the attestation service verifies the integrity of the trusted process by comparing the measurements of its components against the values provided by the function provider’s policy during the initialization phase. Beyond this point, subsequent invocations of the same function become a warm start. Wallet’s attestation service only has to measure the function input and output and can generate a report based on prior measurements of the monitor, the zygote, and the invoked function. When several serverless functions are chained, the extended attestation report includes all their measurements. This approach reduces the attestation time while ensuring verifiability for the entire data chain.
6 Security Analysis
| Attack vector | Mitigation |
| From guest OS | |
| Read the Wallet’s memory | VMPL protection |
| Read user data in the request | Data encryption |
| Load crafted zygotes/functions/inputs | Attestation |
| Load compromised external file | Measurement check |
| From host and hypervisor | |
| Access VM’s memory | CVM protection |
| DMA to the VM’s memory | CVM protection |
| Inject malicious interrupts [138, 139] | Alternate injection [9] |
| Return invalid CPUID values [93] | CPUID pages [14] |
| Launch compromised images | Attestation |
| Manipulate user request | Protocol encryption |
| From co-located functions | |
| Access other trustlet’s memory | Page table protection |
| Reuse-based attacks [167] | Recreate trustlets |
6.1 Attack vectors & Wallet’s mitigations
Table 4 summarizes potential attack vectors and Wallet’s mitigations.
Guest OS. First, we consider attacks originating from the untrusted guest OS within the CVM. An attacker may attempt to access data of a serverless request. However, this data is encrypted with a user key that only the trusted monitor possesses and is inaccessible to external entities, provided that the user secrets are not compromised. On top of that, modifying the VMPL configuration via the RMPADJUST instruction [14] is prohibited in the guest OS (VMPL-2). Thus, any attempt from the guest OS to access memory regions of the trusted monitor and Trusted Processes fails thanks to VMPL protections, which are set by the trusted monitor and trap such accesses.
Further, an attacker could load arbitrary zygotes and functions. However, Wallet’s attestation service blocks this process if there is a measurement discrepancy from the values in the function provider’s policy. Replay attacks by replacing user input are also detected as users can validate the integrity of the input based on the content of the attestation report. Lastly, if an attacker intercepts a file request from a trustlet to the nested filesystem (§ 5.2) and returns a compromised file, the LibOS verifies the file’s measurement against its expected value in Wallet’s manifest, rejecting any tampered files.
Host and hypervisor. CVMs employ mechanisms to protect CVM’s private memory and provide ways to retrieve trusted system information from within the CVM. Precisely, unauthorized access to CVM memory is blocked by AMD SEV-SNP’s reverse map page table (RMP), which prevents the host, hypervisor, and other co-located VMs from accessing CVM’s private memory. Direct memory access (DMA) attacks from the host are also infeasible since DMA is prohibited by hardware. Further, to address potential threats to compromise the guest via arbitrary interrupt or exception injection [138, 139], SEV-SNP introduces alternate interrupt injection [9], restricting the hypervisor to specific exception types. On top of that, an attacker may try to disrupt the system by tampering with the CPUID instruction [93] return values. As a countermeasure, SEV-SNP provides a CPUID-page mechanism, which ensures correct CPUID values, validated by the ASP.
Attempts to compromise Wallet by launching a tampered trusted monitor, bypassing it altogether, or directly booting Wallet in a standard VM are detected during Wallet’s remote attestation workflow. Attestation reports include invalid measurement values if altered software is used. Attempts to deploy Wallet as a normal VM can be identified, as, in this case, Wallet fails to produce a valid report from the ASP. Even crafted reports are detected, since only the CPU vendor’s key, securely managed by the ASP, can sign genuine attestation certificates. Lastly, establishing secure connections (e.g., TLS) protects against network communication manipulation by ensuring the confidentiality and integrity of network packets.
Co-located serverless functions. Wallet prevents attacks from co-located functions. An attacker might create a serverless function that exploits a runtime vulnerability aiming to access memory belonging to another trustlet. However, each trustlet runs in ring3 and has its own page tables managed by the trusted monitor, which prohibits access to other trustlet’s memory. Further, to prevent reuse-based attacks [167], Wallet’s process runtime recreates trustlets when receiving a request from a different user, resetting the potentially compromised residual state due to a previous invocation.
6.2 Formal verification of Protocols
We formally verify Wallet’s differential attestation protocol, as depicted in Figure 7, using the Tamarin Prover [105, 50] under the Dolev-Yao [51] attacker model. The verification considers two key properties for Wallet’s differential attestation protocol: (i) secrecy, secrets (e.g., encrypted data, private keys) remain undisclosed unless explicitly compromised; and (ii) authenticity, guaranteeing that attested results correspond to correct function execution with validated inputs. We model the steps specific to Wallet in detail.
More precisely, an unbounded number of protocol flows are considered simultaneously, with their individual states represented as part of a global multiset. State transitions are encoded as rewriting rules, which are one of the input formats for the Tamarin prover, and long-term secrets are treated as compromisable. Messages are considered atomic, and cryptographic functions (e.g., hashing, encryption) are assumed to be perfect, i.e., without side effects or collisions, and known to the attacker. We further assume the correct functionality of the underlying hardware, i.e., the ASP does not leak its secrets and produces correct reports, and report verification infrastructure. Such external dependencies are treated as a black box.
Our model is based on action facts, which denote events in the protocol trace. We use the Secret(s) action fact to explicitly mark secret information s, in conjunction with the builtin K(x) action fact, which marks an adversary obtaining information x. We further introduce the UsrTrustRes(…) action fact to indicate that a user trusts, after the last protocol step, that the obtained result was indeed computed on a trusted software stack.
Rules. To translate the protocol specification to rules, which operate on a global state, we identify: (i) the necessary inputs from the network, as well as, the persistent states for each Agent for the rules left-hand side, (ii) the resulting outputs on the network, as well as, any modifications to the persistent states of each Agent for the rules right-hand side, (iii) any checks performed by an Agent, which may translate to restrictions of the rule transitions.
This approach enables a systematic derivation of the model from the specification for most of the protocol, some parts, however, require additional consideration:
The preparation phase contains an exchange of the public function key, between the user and the function provider, which is not further specified, for the formal analysis we model this exchange as a secure channel (Figure 8). The motivation for this is, that we do not consider attacks on this information exchange. The attacker is still able to obtain the function’s public key, which we model with
Out(pk(~func_priv)) on the right-hand side of the key generation rule, but unable to provide the user false information in the preparation phase.
rule ChanOut_S:
[ Out_S($A,$B,x) ]
--[ ChanOut_S($A,$B,x) ]->
[ !Sec($A,$B,x) ]
rule ChanIn_S:
[ !Sec($A,$B,x) ]
--[ ChanIn_S($A,$B,x) ]->
[ In_S($A,$B,x) ]
The initialization phase is largely based on Tamarin’s diffie-hellman builtin, for modelling the key exchange. To model the report generation, verification and user-data retrieval as a black-box we introduce the functions gen, verif and getD respectively. The expected behaviour of these is expressed in the following equations, where m refers to the machine, d to the software measurement, and u to the retrievable user-data:
In this phase we also included an unrestricted machine_init to be able to set up an infinite number of machines and thus infinite potential Wallet instances. This, together with the way we modeled the function provider in the previous phase, allows us to model an unbounded number of Wallet instances, machines, functions, providers, and users all while ensuring they are unable to interfere to violate our desired properties.
The remainder of the rules are straightforward translations from the protocol.
Lemmas. We use to denote that action fact a occurred at time . Leveraging these action facts, we verify the following lemmas for Wallet’s differential attestation protocol:
Secrecy lemma: If some data (e.g., function input/output) is declared as secret s, it remains undisclosed to attackers unless one agent A having access to it (KnowSecret(A, s)) is explicitly compromised (Compr(A)).
(3) -
Authenticity lemma: If a user trusts a result, after verifying the differential attestation report, then the result res was computed on the trusted monitor in a valid state tm, and with the expected zygote zy, trustlet tr, and input in, which is modeled by the function.
Tamarin verifies the specified properties by showing that there is no trace that leads to a falsification of these lemmas. Thus, we show that the attestation protocol ensures secrecy, authenticity, and correctness under the specified assumptions.
However, to increase confidence in the correctness of our model we include a simple sanity check, which ensures that it is possible to reach this point in the protocol. For this we use this lemma:
If this lemma is falsified, then the authenticity lemma would trivially hold, since it would be a for-all condition on an empty set. However, this would likely indicate an issue in our model and not the proper functioning of the protocol, therefore we include the above sanity check to rule out this edge case.
| Information | Agents with access |
| User’s Function input | User, TrustedMonitor, Function Provider |
| Output encryption key | User, TrustedMonitor, Function Provider |
| Function’s private key | TrustedMonitor, Function Provider |
For the secrecy analysis, we were also required to explicitly state the agents that can potentially access, i.e., obtain in plain text, certain secret information, as depicted in Table 5. We found the sets of agents to be minimal, as further reduction caused the secrecy lemma to be violated. Notably, this required us to include the Function Provider as an agent, with potential access to the User’s function input and output. The reason for this is the private function key, which is known to the function provider.
The persistence of these long-term keys and their use also prevented us from verifying a stronger secrecy property called perfect forward secrecy. This would ensure that information will remain secret, even if the attacker is able to record the communication and compromize some long-term key in the future. Checking for this property entails making a small modification to the secrecy lemma:
For our protocol, Tamarin is able to correctly identify that this lemma is violated since compromising the long-term private key of a function enables the decryption of any previous request to that function. We consider this attack vector out of scope for this paper, as it relies on compromising a key component of our infrastructure. Future work may explore extensions to the protocol, that address this, to better contain an initial compromise.
7 Implementation
Trusted monitor. We implement our prototype of Wallet for Linux environments. We base Wallet’s trusted monitor on COCONUT-SVSM [38]. COCONUT-SVSM is a service module that operates at VMPL-0 and aims to provide services (e.g., vTPM [114]) to a guest running at VMPL-2. We extend COCONUT-SVSM to support trusted processes in VMPL-1 and incorporate Wallet’s differential attestation service. We base our LibOS for trusted processes on Gramine [27, 39] and implement a backend to enable its execution at VMPL-1. Wallet’s current prototype uses a patched Linux version v6.8 (host OS) and v6.5 (guest OS), configured to support COCONUT-SVSM.
Communication protocol. The COCONUT-SVSM implements a guest communication protocol defined by AMD [12], enabling the guest OS to request services from the COCONUT-SVSM using the VMGEXIT instruction. We extend this protocol to support the monitor system calls (Table 2) by adding a new service type and implementing its corresponding handlers.
For communication between the trusted monitor and trustlets, we enable #VC (VMM Communication Exception) reflection [14], allowing the trusted monitor to trap #VC during trustlet execution. We define a communication protocol based on the cpuid instruction, leveraging the unused hypervisor CPUID Leaf range [13]. Precisely, trustlets set a service type in the RAX register and execute the cpuid instruction, triggering #VC. In turn, the trusted monitor inspects the RAX value, executes the corresponding monitor service, if applicable, or the normal cpuid instruction, and returns the result.
LibOS. Our LibOS for trusted processes uses Gramine [27, 39] with a backend for VMPL-1 execution. Gramine defines a Platform Adaptation Layer (PAL) host ABI [40] to delegate specific operations to the host. We implement our PAL as a shim layer for our trusted monitor, which allows Wallet to run the Python runtime as a trusted process. Its core functionalities include memory allocation and external file access, and it leverages Wallet’s communication protocol to perform its PAL ABI operations requests.
For memory allocations, the trusted monitor handles the requests from the PAL. Specifically, it allocates memory, adjusts the VMPL access permissions using the RMPADJUST instruction, and updates the trustlet’s page table entries. For external file access, the trusted monitor returns a special value to the guest OS to request the desired file. Then, the guest OS reads the file and invokes the invoketrustlet() monitor call. Subsequently, the trusted monitor copies the file data into the trustlet’s memory and resumes its execution. The trustlet validates the file integrity before using it by comparing its measurement with a specified value in the manifest. For the memory backend, the trusted monitor secures the memory pool and validates each page during its boot to optimize memory allocation at runtime. When a trustlet requests memory, the trusted monitor allocates memory from the pool, and updates the trustlet’s page table entries accordingly.
Copy-on-write handling. Gramine LibOS operates in ring3 and does not contain exception handling code. To realize copy-on-write-based forking, we develop a minimal exception handler that executes in ring0 within a trustlet (VMPL-1). The trusted monitor configures the GDT (Global Descriptor Table), IDT (Interrupt Descriptor Table), and TSS (Task State Segment) of the trustlet to ensure that it uses the handler upon launch. On a page fault, the page fault handler delegates the actual handling to the trusted monitor via a monitor call. The trusted monitor resolves it by updating the page table entries appropriately and then resumes the execution of the trustlet.
Attestation service. Wallet leverages COCONUT-SVSM’s get_regular_report() to retrieve a signed AMD SEV-SNP report and, then, stores it securely in the trusted monitor’s memory. To calculate the measurements of individual trusted processes components (e.g., zygote image, trustlet), Wallet uses the SHA-512 algorithm (same as the SEV-SNP’s measurement algorithm [14]). These measurements are cached in the zygote and trustlet contexts for faster retrieval. The monitor invokes the attestation functions internally when fetching a zygote or creating a trustlet, validating them against the function provider’s policy by comparing the calculated with the provided measurements. Finally, on a function invocation, the measurements of the used software components and the function input and output are appended to the trusted monitor’s attestation report, which is then signed using the private key provided by the function provider and forwarded to the user for verification of the entire execution.
Guest OS components. Wallet includes a kernel module that allows a guest application to make monitor calls. In addition, we implement a Python library that interacts with the trusted monitor through this kernel module. We also extend the SeBS benchmark suite [34] to support running serverless functions with Wallet using this Python library.
8 Evaluation
We evaluate Wallet by analyzing its end-to-end performance (§ 8.2),the performance of its operations (§ 8.3), its resource efficiency (§ 8.4), its communication network performance (§ 8.5), and its scale-out capabilities (§ 8.6).
8.1 Experimental Setup
Testbed. We perform our experiments on an AMD SEV-SNP-enabled server with an AMD EPYC 7713P CPU (64 cores, hyperthreading disabled) and 1024 GB of DDR4 DRAM (1664 GB/DIMM). The server runs NixOS 24.11 with an AMD SEV-SNP-enabled Linux kernel (v6.8.0). Each VM uses an Ubuntu-22.04 with a VMPL-enabled Linux kernel (v6.5.0).
Variants. We use the baselines summarized below.
| Variant | Execution environment |
| Native | Bare-metal instance on Linux |
| LibOS (Gramine) | Gramine LibOS on Linux |
| Container (Kata) | Kata containers runtime with QEMU/KVM |
| VM (KVM-Linux) | Standard VM with Linux guest OS |
| CVM (SEV-SNP) | AMD SEV-SNP VM with Linux guest OS |
| Wallet | Our Wallet system |
Workloads. For performance evaluation, we use Python functions from the SeBS benchmark suite [34]. The current Wallet prototype lacks external network and fork() support; therefore, we exclude “uploader”, “crud-api” (requires external networking), and “video-processing” (requires running ffmpeg as a standalone process). Additionally, we conduct a simulation-based study with Azure Functions production traces [165, 142].
8.2 End-to-end Performance
Methodology. We run the SeBS server in each variant and co-locate the client on the same host. Wallet does not include the time of measurement calculation for a fair comparison with the other baselines. We run each experiment five times and report the average.
End-to-end latency. Figure 9 presents the end-to-end client latency of each SeBS function for each variant. In the cold start case, Wallet (Lukewarm) achieves 93.44% speedup over CVM (SEV-SNP) on average, and is 83.99%, 51.04%, 32.85% faster than VM (KVM-Linux), Containers (Kata), and LibOS (Gramine). Even in the “cold”est case (i.e., no loaded zygote), Wallet is on average 85.10% faster than the CVM (SEV-SNP) baseline and 63.62% faster than the VM. When compared to the Containers (Kata), and LibOS (Gramine), Wallet incurs 11.25% and 52.57% performance overhead, respectively. In the warm start case, the performance difference among variants becomes smaller as the existing function environment is reused. Still, Wallet is 14.97%, 7.53% and 1.04% faster than the CVM, VM and Containers (Kata) case respectively while it shows a 15.63% performance overhead relative the LibOS (Gramine). Wallet’s performance gains mainly stem from its efficient function startup, run-to-completion execution model that minimizes VMEXITs, and in-memory filesystem. Wallet exhibits higher latency if the input/output size is large due to data transfer overhead between different VMPL levels (e.g., image-recognition has 100MB input) (§ 8.3).
Invocation latency. Figure 10 shows the cumulative distribution function (CDF) of invocation latency under cold starts, which primarily affect the invocation latency. Wallet (Lukewarm) has p50/p99 (50th and 99th percentile) latencies of 0.019/1.06 s. It achieves ms-scale p50 startup latency, comparable to the latest (non-confidential) fork-based serverless systems [26, 53, 116, 52, 96]. On the other hand, Wallet (cold) has 1.57/12.34 s latency, which is at the levels of Containers (1.93/2.10 s) and much lower than the CVM (21.25/22.10 s) and the VM (7.5/8.0 s) variants. It also exhibits a larger standard deviation (3.51 s), which originates from the time needed to load the zygote and the variably-sized functions (§ 8.3).
8.3 Performance Analysis
Boot time analysis. Figure 1 presents the boot time breakdown. For VM-based variants, the majority of the boot time is consumed by VMM (QEMU) initialization and guest OS startup. The CVM incurs additional overhead due to extra management tasks (e.g., measurement of initial state, memory validation) [112]. Although Wallet’s environment setup takes 4.05 s longer than the CVM (without memory preallocation (§ 8.3.1)) because of the added setup time for the monitor, Wallet’s cold start only takes 2.38 s by eliminating the CVM initialization time. Wallet’s lukewarm start further optimizes the boot time by employing a fork-based start (10.3 ms).
Runtime processing cost. Further, we analyze the runtime processing of Wallet, focusing on zygote and trustlet creation, and input/output transfer. Each process involves transferring data between the guest OS and the trusted monitor, followed by updating the VMPL level and the trustlet’s page tables. As Figure 10 illustrates, the zygote creation is responsible for most of the runtime initialization. The current prototype bundles required libraries in the base image (e.g., numpy, pytorch), increasing the zygote size, ranging from 60 MB to 691 MB (image-recognition) in SeBS. However, zygote creation occurs only during cold starts, and the CoW mechanism enables efficient memory sharing among trustlets using the same zygote image (§ 8.4).
On the other hand, the trustlet creation and data transfers’ duration depends on the function and I/O size. Each SeBS function is less than 4 KB. Thus, the trustlet creation requires ms. However, larger data transfers, such as in dna-visualisation (112 MB output), take ms. If the data fits within a 4K page, the copying time is ms (e.g., graph-mst and graph-pagerank).
Measurement cost. Figure 2 shows the breakdown of the time taken to calculate the SHA512 measurement of an empty Python function. The calculation time is proportional to the data size. In the cold start case, the measurements include those of the trusted monitor, as well as the zygote (60 MB), which takes around 1.1 s. The current prototype does not support CPU acceleration for SHA, justifying the lower performance. However, for the lukewarm start, Wallet only needs to measure the function, input, and output (typically less than 4KB), dropping the calculation time below ms.
8.3.1 Effectiveness of Optimizations
Memory preallocation. Memory validation [14] is expensive as it involves VMEXIT for state updates. In our setup, validating a 4K page takes 24 µs (6ms/MB). Although this cost is an one-time overhead, it affects the initial creation time of zygotes and trustlets. Memory preallocation removes this validation cost at runtime in exchange for an increased bootup time. For the SeBS benchmark, the preallocation improves the performance by 49.22% in the cold start and 4.72% in the warm start case on average, over the non-preallocation version. However, preallocating 16 GB of memory increases the boot time by 238 s, highlighting the trade-off.
Copy-on-Write (CoW). We further investigate the impact of CoW. Performance-wise, CoW is crucial for achieving rapid trustlet creation and, thus, fast lukewarm starts (Figure 10). CoW reduces the average trustlet creation time of SeBS functions from 66 ms to 0.11 ms. However, it incurs additional overhead during execution due to page faults. This overhead becomes noticeable mostly in the warm starts, but still remains almost negligible (1%), promoting CoW’s adoption.
8.4 Resource Efficiency
Memory footprint. We evaluate the memory footprint of Wallet while running SeBS functions. Figure 10 presents the result. “CoW shared” represents the memory shared with CoW among trustlets and their base zygote, whereas “Non-shared” denotes the memory exclusively allocated for a trustlet during execution. Although the amount of allocated memory (“Non-shared”) varies depending on the workload, on average, trustlets share 170 MB of memory with the base zygote, reducing the total system memory consumption.
Function density. Memory sharing enables Wallet to pack more functions per node. We analyze the memory consumption for increasing numbers of concurrently-running empty Python functions, shown in Figure 2. We enable kernel same-page merging (KSM) [45] on the host to de-duplicate identical memory pages. For Wallet, the total memory is calculated as the base zygote shared memory (147 MB) plus 60 KB of non-CoW-shared memory per function. As KSM cannot deduplicate encrypted memory, CVM memory consumption is proportional to the number of functions, leading to a higher per-function memory overhead. Further, the number of CVMs in a single node is limited to the number of encryption keys (509 in our environment). In contrast, Wallet achieves significantly lower memory usage, realizing high function density.
8.5 Communication Analysis
Methodology. To evaluate Wallet’s inter-function communication, we measure the communication time of invoking multiple functions in a chain by varying the data size and the number of functions in the chain. All functions are deployed on the same host. In Wallet, each function creates its communication endpoint for input and output, while other baselines use TCP networking.
Communication latency. First, we show the communication latency between two functions under various data sizes. Figure 1 shows the result. This result indicates that Wallet achieves lower latency than CVMs (1.4-27), as well as traditional VMs and Containers (Kata), across message sizes, thanks to the zero-copy-based data objects.
Latencies of function chains. Next, we show the latencies of function chains. As shown in Figure 11, Wallet outperforms all baselines across all chain lengths, achieving 16.7-30.2, 11.8-20.9, and 15.2-30.1 lower communication latency than CVMs, Containers, and traditional VMs, respectively. Wallet’s performance advantage grows with chain length, demonstrating the efficiency of its data-centric I/O for inter-function communication compared to TCP-based networking used by other approaches, highlighting Wallet’s suitability for complex workloads with function chaining.
8.6 Scale-out Performance
Methodology. We develop a simulator modeling real-world serverless deployments with a central scheduler and multiple nodes, each having fixed execution slots and LRU caches for warm functions. The scheduler processes trace data chronologically, prioritizing nodes already caching the requested function. If none are available, it assigns requests to free nodes or queues them. Function execution duration gets increased based on its boot type (cold/lukewarm/warm), using timings sampled from our microbenchmarks (Figure 1). The lukewarm boot, unique to Wallet, occurs when a node caches another function from the same application, simulating zygote sharing. We use Azure Functions production traces [165, 142] sampled with InVitro [154], comprising 4k functions with 4.1 million invocations over 30 minutes.
Invocation latency and per-function slowdown. We run the simulator with 100 nodes, each having 32 execution slots and a cache size of 32. Figure 12 shows the CDF of invocation latency: at p50/p99, CVM and VM experience delays of 489/881 s and 1.3/33 s, respectively, while Wallet achieves just 5 ms/1.5 s. Figure 12 reports the per-function slowdown, which increases as function execution time decreases and invocation latency dominates. At p50/p99, CVM slowdowns reach 2,682/391,257, VM slowdowns are 13.98/8,658, while, in contrast, Wallet’s slowdowns are merely 1.02/7.06.
Node scalability. We also evaluate the invocation latency for various numbers of nodes. Figure 12 shows that, compared to the other baselines, Wallet achieves lower latency with a smaller number of nodes. For example, Wallet has 5 ms p99 latency with 100 nodes, while, even with 150 nodes, the CVM’s p99 latency is 50 s, and the VM’s is 12 s. These results highlight Wallet’s effectiveness in a large-scale environment.
9 Related Work
Serverless computing. Serverless computing has transformed cloud deployment. Open-source frameworks (OpenWhisk [48], OpenLambda [63], OpenFaaS [47]) and major cloud providers (AWS Lambda [141], Azure Functions [108], Google Cloud Functions [60]) offer widely used architectures. Recent research [80, 132, 81, 53, 165, 128, 146, 5, 116, 144, 26] mainly focuses on improving serverless isolation, performance and efficiency, while Wallet targets the security challenges and leverages TEEs to form a lightweight confidential computing system for secure serverless deployments.
Confidential computing. Confidential computing, based on hardware TEEs [16, 9, 77, 17, 18, 76, 90], is widely adopted for protecting data and code in untrusted clouds. TEE technologies can be broadly categorized into process-based (e.g., Intel SGX [76]) and VM-based TEEs (e.g., AMD SEV-SNP [9], Intel TDX [77], ARM CCA [17]). Extensive recent research [118, 61, 167, 4, 11, 88, 56, 159, 129, 65, 58, 113, 134, 97, 155, 160, 92, 157, 163, 145, 57, 21] aims to minimize TCB, and provide intra-VM isolation, secure data at rest, among others.
Specifically, Veil [4], and NestedSGX [157] deploy secure services in CVMs using SEV-SNP VMPL. Erebor [163] adopts intra-kernel privilege isolation to provide sandboxing in CVMs. Unlike these, Wallet targets serverless deployments, and leverages CVM partitioning for minimal TCB, fast boot, and efficient function communication via shared memory.
Confidential serverless computing. Many research works aim for secure serverless computing architectures using confidential computing [123, 7, 94, 140, 84, 25, 55, 152, 119]. ServerlessCoCo [140] comprehensively analyzes overheads in CVM-based serverless deployments. Plugin Enclaves [94] proposes a hardware-based approach with Intel SGX for efficient confidential serverless computing. Cryonics [84] reduces startup times using snapshot-based SGX enclaves. In contrast, Wallet materializes confidential serverless computing by providing a CVM-based lightweight system with easily attestable and deployable trusted processes.
Concurrently with our work, COFUNC [143] proposes a library OS-based confidential containers tailored for serverless functions. COFUNC employs a microkernel architecture, running multiple containers with the same CVM while enforcing isolation. Contrary to Wallet, COFUNC offloads container management tasks to the serverless runtime that runs on the host. On the other hand, Wallet allows for the reuse of the existing OS and serverless framework within the CVM, leveraging its nested confidential execution architecture.
Lightweight virtualization. Lightweight virtualization techniques [3, 42, 87, 27, 39, 88, 85, 102, 100] aim to reduce the VM overheads while maintaining isolation. Firecracker [3] introduces microVMs for serverless workloads. Kata Containers [42] deploy lightweight VMs for container applications, whereas Unikraft [87] creates specialized unikernels for applications. Gramine LibOS [27, 39, 88] runs unmodified apps in SGX enclaves and TDX VMs [88]. On the other hand, Wallet leverages these concepts with a minimal Gramine LibOS instance for each of its trusted processes, enabling low boot and attestation times, and seamless application deployment. Unlike Gramine-TDX, Wallet’s LibOS is only running in a certain privilege level within the CVM and does not have system-wide privileges. Further, in contrast to Gramine’s single-tenant model with its existing backends, Wallet provides a framework for managing multiple instances within the same CVM, enabling multitenancy and optimized inter-function communication via data objects among trusted processes while maintaining function instance protection.
Networking/IO for confidential computing. TEE networking is typically slow due to the added overhead of encryption and data copying with multiple context switches [92, 112, 158] Several works optimize their processing overhead with polling at the cost of increased CPU usage [150, 151, 20, 21]. Compared to them, Wallet achieves efficient inter-function communication through its data-centric shared-memory-based I/O architecture while ensuring isolation.
10 Conclusion
In this paper, we present Wallet, a lightweight confidential serverless computing system. Wallet leverages intra-CVM partitioning mechanisms and efficiently consolidates security-critical functionalities into a compact, privileged trusted monitor, resulting in a small TCB. Through its zygote mechanism, Wallet achieves fast boot times while optimizing memory management and reducing the communication cost between serverless functions via its data-centric I/O architecture. Further, it provides a formally verified differential attestation mechanism to ensure end-to-end trust with reduced attestation times. Overall, Wallet introduces low-performance overheads while exposing a minimal attack surface, making it ideal for security- and latency-sensitive serverless deployments.
Artifact. Wallet is publicly available at https://github.com/TUM-DSE/Wallet-VMPL along with its entire experimental setup.
Acknowledgments
We thank our shepherd, Prof. Anurag Khandelwal, and the anonymous reviewers for their helpful comments. We also thank Prof. Nuno Santos for valuable discussion and feedback on this work.
This work was supported in part by an ERC Starting Grant (ID: 101077577) and the Chips Joint Undertaking (JU), European Union (EU) HORIZON-JU-IA, under grant agreement No. 101140087 (SMARTY), the Intel Trustworthy Data Center of the Future (TDCoF), and Google Research Grants. The authors acknowledge the financial support by the Federal Ministry of Research, Technology and Space of Germany in the programme of “Souverän. Digital. Vernetzt.”. Joint project 6G-life, project identification number: 16KISK002.